Circular No. 7687

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2001 Q2 Additional astrometry by P. Pravec and P. Kušnirak with the Ondřejov

0.65-m f/3.6 reflector + CCD:

2001 UT  α_{2000}  δ_{2000}
Aug. 20.07522  5^{h}41^{m}32.^{s}03  +27^{o}34′34′′.1
20.08395  5 41 35.00  +27 34 23.6

Visual m_{1} estimates: Aug. 20.09 UT, 10.4 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 20.31, 10.8 (P. M. Raymundo, Salvador, Brazil, 0.15-m refractor).

Corrigenda. On IAUC 7686, lines 2 and 12, for Cyprus Hills, read Cypress Hills Interprovincial Park,

NOVA CYGNI 2001 No. 2 S. Nakano, Sumoto, Japan, reports the independent discovery of this

nova (cf. IAUC 7686) by K. Hatayama (Niigata, Japan) on two 1-min exposures taken on Aug. 18.625 UT with a 100-mm f/5.6 camera lens and Fuji Superia 400 film that show the nova at mag 7.0 (limiting mag 10). Nothing was visible at this position on Hatayama’s films taken on 2000 Sept. 29, Oct. 22, 2001 July 21, and July 25. T. Kato, Kyoto University, reports the following limiting magnitudes (nova not seen) from unfiltered CCD images: Aug. 2.654, [13.2 (A. Takao, Kitakyushu, Japan); 11.520, [14.3 (Y. Sano, Nayoro, Japan); 12.699, [14.1 (Takao); 17.628, [14.1 (Takao).

CCD photometry by G. Sostero and A. Lepardo, Remanzacco, Italy (0.3-m Baker camera): Aug. 19.84 UT, V = 6.66, B − V = +1.1, V − R_{c} = +0.5, R_{c} − I_{c} = +0.9. CCD photometry by J. D. West, Mulvane, KS (0.25-m refl.): Aug. 20.098, V = 6.89; 20.138, V = 7.08; 20.087, I_{c} = 4.48.

CI AQUILAE B. E. Schaefer, University of Texas at Austin, writes: “The recurrent

nova CI Aql has now faded back to near its normal quiescent state with an eclipsing light curve varying from V = 15.3 to 16.2, as seen from Aug. 4.1 to 10.3 UT with the 2.1-m telescope at McDonald Observatory. This implies that the system is no longer a supersoft x-ray source and that the hydrogen content of the white dwarf envelope is X ∼ 0.4 (Hachisu

and Kato, http://xxx.lanl.gov/abs/astro-ph/0104461). The formerly

prominent secondary minimum is no longer visible; instead, the magnitude fades linearly by 0.10 mag between phases 0.35 and 0.65.”

2001 August 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green